# Inducing Group Fairness in LLM-Based Decisions

### v1

@@ Abstract
Prompting Large Language Models (LLMs) has created new and interesting means for classifying textual data. While evaluating and remediating group fairness is a well-studied problem in classifier fairness literature, some classical approaches (e.g., regularization) do not carry over, and some new opportunities arise (e.g., prompt-based remediation). We measure fairness of LLM-based classifiers on a toxicity classification task, and empirically show that prompt-based classifiers may lead to unfair decisions. We introduce several remediation techniques and benchmark their fairness and performance trade-offs. We hope our work encourages more research on group fairness in LLM-based classifiers.
Inducing Group Fairness in LLM-Based Decisions
James Atwood     Preethi Lahoti     Ananth Balashankar     Flavien Prost     Ahmad Beirami Google DeepMind
@@ 1 Introduction
Large Language Models (LLMs) have shown impressive performance across tasks and are now being deployed across many high-stakes applications such as financial Wu et al. (2023) or medical Singhal et al. (2023) domains. In particular, zero-shot LLM-based classifiers Wei et al. (2022a); Anil et al. (2023) have been shown to achieve state-of-the-art performance on several natural language classification benchmarks and are being widely adopted for decision making. More recently, such classifiers are even used as a reward signal to align the model to AI Feedback Bai et al. (2022). Hence, it is important to investigate the fairness of such classifiers to different demographics.
While LLMs have been broadly studied in the generative case for diversity, stereotypes, gender bias, and toxicity Nadeem et al. (2021); Liang et al. (2021); Deshpande et al. (2023); Lahoti et al. (2023), their fairness in classification problems remains under-explored. In this paper, we investigate specifically if zero-shot and few-shot classifiers satisfy group fairness notions such as Equality of Opportunity (EO) Hardt et al. (2016), measured as the difference of false positive rates (FPR) between demographic groups.
Not surprisingly, we find that the prompted zero-shot LLM-based classifiers demonstrate a significant gap in FPR across multiple demographic groups, with Muslim and Jewish groups having 89% and 48% higher FPR as compared to the Christian group in the Civil Comments toxicity detection benchmark  Borkan et al. (2019). The gap is further increased when we compare few-shot prompting based LLM classifiers with Muslim and Jewish groups having 124% and 71% more FPR compared to the majority group.
Fairness mitigation for LLM-based classifiers is challenging due to lack of access to pretraining data and procedures. In addition, little to no fine tuning is usually done in zero-shot and few-shot classification tasks, which makes it hard to remediate through standard fairness remediation techniques for classifier fairness Hort et al. (2022).
To mitigate this gap, we benchmark the effectiveness of prompting as well as regularization-based remediation methods. In the prompting methods, we rely on instruction-following Kojima et al. (2022) and study the effectiveness of group agnostic and group aware prompts. We find that these prompt-based methods are unable to decrease the FPR gap - with Muslim and Jewish groups still having FPR about 40% higher than the majority Christian group. Then we study two in-processing Prost et al. (2019); Beutel et al. (2019) and post-processing Ţifrea et al. (2024) remediation methods by learning a softmax classification layer on the final embeddings of the LLMs, which achieve better fairness-performance trade-offs. Note that the in-processing baseline is only applicable to the fine-tuned model, and cannot be applied in the zero/few-shot classification with no training.
@@
  • We benchmark the group fairness of LLM-based classifiers and show that they do not satisfy Equality of Opportunity along identity aspects such as religion, race, ethnicity, sex.
  • We introduce three remediation techniques: prompt-based, in-processing, and post-processing. We highlight that prompting-based remediations which rely on instruction tuning cannot achieve lower false positive rates, and that methods that operate on the embeddings of the LLMs (in/post-processing) are needed to achieve better fairness-performance trade-offs.
@@
  • Our findings suggest that prompt-based methods are not effective for group fairness remediation. Additionally, in-processing remediation achieves better fairness-performance trade-offs than post-processing methods.
Related Work.
@@
Tamkin et al. (2023) provided a method for evaluating how biased a language model may be by generating hypothetical prompts with group information. This was then used to make decisions by fitting a mixed effects model and quantified the degree of bias. The authors of the Flan-T5 model Chung et al. (2022) published group-level performance of a toxicity classifier fit to the Civil Comments Identity Borkan et al. (2019) dataset. Baldini et al. (2021) explored remediation methods for achieving equalized odds for different embedding-based classifier models. To our knowledge, this is the first paper that proposes and empirically evaluates methods for remediating zero-shot and few-shot classifiers drawn from LLMs with respect to equality of opportunity fairness. The related work for fairness in classical classification can be found in Appendix A.
@@ 2 Problem Setup
Throughout this paper, we use PaLM 2 S Anil et al. (2023) as the base model. We also use PaLM 2 L model in a model transfer experiment. We explore two settings: a ‘zero-shot’ setting, and a ‘fine-tuned’ setting where we improve the classifier performance by training a classifier, or head, on top of the final representation of the model.
For both the zero-shot and fine-tuned case we begin by wrapping the text snippet to be classified in ‘wrapper text’ that encourages the LLM to make a decision. As an example, consider a post with the text ‘first post!’ In this toy example, in order to encourage the LLM to make a decision, we wrap the text to obtain ‘is “first post!” toxic? please respond with Yes or No.’ We then send this to an LLM and ask for either a text answer (zero-shot) or model activations (fine-tuning) from the LLM.
Zero-Shot Classifier. We use a scoring-based zero-shot method for adapting LLMs to classification tasks. A schematic diagram is given in Fig. 2(a) in the appendix. Our goal is to take a text post and a classification target as input and return a distribution over negative and positive classes as output. To accomplish this, we send wrapped text to an LLM and ask it to score the ‘Yes’ and ‘No’ tokens to obtain positive and negative class scores. We then apply a softmax to obtain a probability distribution and use that distribution when making decisions. This is similar to the methodology that is followed by Chung et al. (2022).
Fine-Tuned Classifier. In the fine-tuned case, we use the activations of the final layer as input, and train a logistic regression model to predict the ground truth labels. The logistic regression model is then used for decision making. This approach can be thought of as a form of parameter efficient supervised fine tuning. A schematic diagram is given in Fig. 2(b) in the appendix.
@@ 3 Experiment Setup
Dataset and Tasks.
We use the Civil Comments Identity Borkan et al. (2019) dataset as it provides the information necessary to evaluate and remediate both group fairness and performance. For processing details, please see Appendix D.
Metrics.
@@
Group fairness remediation baselines.
We now evaluate several remediation methods. We propose a family of prompt-based methods that ask the model to be fair at different levels of specificity (in general, to demographic groups, and to a specific group). We also include in-processing and post-processing methods that are adapted from successful classical approaches:
  • Post-processing. We apply FRAPPE with an MMD regularizer base (Ţifrea et al., 2024).
  • In-processing. We apply the in-processing MMD regularization of Prost et al. (2019) to model finetuning. This baseline is inapplicable to zero/few-shot prompt-based classifiers.
  • Prompt PBF. ‘Please be as fair as possible when making a decision’ is appended to the prompt.
  • Prompt PBF2SG. ‘Please be as fair as possible when making a decision about comments about religious groups or that mention religion’ is appended to the prompt.
  • Prompt PBF2TG. ‘Please be as fair as possible when making a decision about comments that mention Judaism or Jewish people’ is appended to the prompt.
For more details about the remediation methods, please see Appendix E.
4 Experiment Results
Benchmarking group fairness without remediation.
First, we evaluate PaLM 2 S model Anil et al. (2023) with respect to equality of opportunity. These experiments follow the classification methodology described in Sec. 2 and use the Civil Comments Identity Borkan et al. (2019) dataset.
Group Zero-shot FPR Ratio Fine-tuned FPR Ratio
Muslim 1.89 2.24
Jewish 1.48 1.71
Table 1: False positive rate ratios that quantify the magnitude of violation of equality of opportunity for a classifier built on PaLM 2 S. We only include the two groups with the highest gaps. Here we used ‘Christian’ as the majority group.
The results of this evaluation for the two groups with the highest ratio gaps are given in Tab. 1 and the full table is given in Tab. 2 (in the appendix). These elevated ratios imply that there is headroom for group fairness remediation. In the rest of the experiments, we will describe remediation approaches and present their empirical performance.
Comparing the Pareto frontiers of group fairness remediation.
(a) Zero-Shot Pareto
(b) Fine-Tuned Pareto
(c) Model Transfer Pareto
Figure 1: A comparison of the Pareto frontiers of in-processing and post-processing techniques. The left plot shows the performance and fairness of zero-shot remediation and the middle plot shows fine-tuned remediation; top plots give results for the Jewish group and bottom plots gives results for the Muslim group. Prompt-based methods are single points indicated by large symbols. Note that the in-processing baseline is inapplicable in the zero-shot case. Each point for in-processing and post-processing is generated by setting different values for $\lambda$ in Equations (2) and (4) in the appendix. The bold point gives the unremediated ($\lambda=0$) case for in-processing and post-processing. The dashed and solid lines give the Pareto frontier where performance can only be gained by sacrificing fairness, for post-processing and in-processing, respectively. The right plot gives the effect of model transfer. We fit a post-processing remediation model to the PaLM 2 S model then compare the effects of applying it to PaLM 2 S (native) versus the larger PaLM 2 L model (transfer). The lines give the Pareto frontier (solid for native and dashed for transfer).
Next, we benchmark the different remediation techniques for performance vs group fairness trade-offs. The results of this are shown in Fig. 1(a). For in-processing and post-processing methods, each point in the plot is generated by varying the regularizer strength (see Appendices E.2 and E.3, respectively for more details). We observe that both methods improve fairness without significantly degrading the performance of the classifier.
We also use the activations of the final layer of the LLM to fine-tune a classifier as described in Sec. 2. The in-processed technique remediates during fine-tuning; the post-processed technique performs a separate remediation step after fine-tuning is complete. Note that this fine-tuned model generally performs better than drawing predictions directly from the LLM as described in Sec. 2. However, it requires access to a supervised dataset for fine-tuning. We compare the Pareto frontiers of fairness and performance for each remediation technique in Fig. 1(b). As before, each point in the plot is generated by varying the regularizer strength, the term that trades off between performance and fairness terms in each technique’s loss function.
There are a few takeaways from these experiments. First, the ability to fine-tune on LLM-derived activations improves the performance of the classifier as compared to zero-shot case. Second, as before, we are able to improve performance without significantly degrading the performance of the fine-tuned classifier. Third, the in-processing technique generally performs better than the post-processing technique in this environment. Finally, prompt-based methods can offer some fairness and performance benefit but are generally less general and less effective than in-processing and post-processing methods.
Transfer of remediation to an unseen model.
This experiment has a different setup than those described earlier in this section. Here, we no longer make the assumption that the LLM has ‘introspective’ abilities and thus can not return its activations to use as an embedding. Instead, we use a ‘third party’ model to embed the prefix of our query and use these embeddings for any task where they are required. More specifically, we use the Google News 128-dimensional embedding model Bengio et al. (2000). [^1] https://www.kaggle.com/models/google/nnlm.
This enables us to operate in environments where drawing embeddings from the LLM is not possible. One interesting case is where we train a post-processing remediation model on one LLM then apply to another; can we reuse the existing fairness model to improve fairness with the new model? Fig. 1(c) gives the results this model transfer learning scenario. Importantly, we find that post-processing model is still able to improve fairness when transferred (although this comes at the cost of higher performance degradation than when applied to the model that it was trained on). This suggests that we can save on inference costs by training our fairness model on a LLM with fewer parameters then apply it to a larger LLM at inference time.
Note also that the Pareto frontier achieved by the smaller PaLM 2 S model in Fig. 1(c), where ‘third party’ embeddings are used, is just slightly degraded from the Pareto frontier given in the upper plot of Fig. 1(a) where model activiations are used as embeddings. This suggests that using ‘third party’ embeddings is a reasonable choice for remediation if model activations are not available.
5 Concluding Remarks
We study the fairness of LLM-based classifiers. We identify that LLM-based classifiers may exhibit group unfairness. We introduce three remediation techniques to improve fairness while maintaining acceptable performance for LLM-based classifiers. We find that prompt-based techniques offer limited benefit and are outperformed by in-processing and post-processing techniques. We also find that the in-processing technique consistently provides favorable performance vs fairness trade-off in the fine-tuned settings.
Based on our results, if remediating a fine-tuned model, in-processing seems to be a more robust approach because it consistently provides favorable fairness versus performance tradeoffs. In other LLM-based classification settings where in-processing cannot be applied (such as zero-shot remediation and transfer tasks) the post-processing technique seems to a promising approach.
We generally find that the prompt-based remediation methods have little to no impact of prompts on fairness, while counter-intuitively, we observe that fairness-oriented prompts may slightly improve performance in some cases for the less specific ‘Please be Fair’ (PBF) and ‘Please be Fair to Super Group’ (PBF2SG) methods. This is not very surprising given that fairness is a distributional issue, and hence simple prompting may not necessarily provide the distribution matching effects that we expect from remediation.
Limitations
  • We find that prompt-based methods are less flexible and effective than in-processing and post-processing methods. However, we do not make an exhaustive search of possible prompts and other researchers may find prompt-based methods that work. Furthermore, it has been observed that LLM capabilities improve as LLM size increases Wei et al. (2022b), and this could have a beneficial effect on prompt-based method effectiveness as LLMs become larger and more capable.
  • Our experiments are focused on equality of opportunity (group) fairness. There is no guarantee that they will generalize to other notions of fairness, and, importantly, their application does not imply that a classifier is abstractly ‘fair.’
  • LLM-based classifier inference is very expensive when compared to simpler models, and the performance of LLM-based classifiers does not yet justify that cost (for example, our LLM-based classifiers are less effective than baseline methods given by the authors of the Civil Comments dataset paper Borkan et al. (2019)). An implicit assumption of this work is that the performance of LLM-based classifiers will improve enough over time to justify their high inference costs and become deployed systems where fairness considerations are in play.
  • We consider only one LLM (PaLM 2) and one dataset (Civil Comments Identity) in English. So it remains to be seen how much our findings generalize. Having said that, given that we already find performance disparities across subgroups in this case, the need for developing fairness remediation techniques is justifiably real.
  • We only experiment with a few-handcrafted prompts for classification, and did not compare against chain-of-thought Wei et al. (2022c), self-consistency Wang et al. (2022), and automated prompt generation Gao et al. (2021) techniques - as adapting them to induce group fairness was not trivial and is left for future work.
  • We do not compare against low-rank adaptation Hu et al. (2021), prompt-tuning Lester et al. (2021), and other parameter-efficient fine-tuning techniques Liu et al. (2022) for the in-processing method.
Appendix A Related Work in Classical Classification Fairness
There is rich literature on fairness in ‘classical’ (i.e. non-LLM-based) classification. In this paper we focus on the ‘equality of opportunity’ notion of group fairness Hardt et al. (2016) which is achieved for a group and classifier when the false positive rate (or false negative rate) of the classifier is the same for instances drawn from that group when compared with instances drawn from the majority group.
Methods for improving group fairness can generally be categorized in three main classes: pre-processing, in-processing, andpost-processing methods. Pre-processing algorithms (Feldman et al., 2015; Zemel et al., 2013; Calmon et al., 2017) transform the biased data features to a new space in which the labels and sensitive attributes are statistically independent. In-processing methods (Kamiran et al., 2010; Ristanoski et al., 2013; Quadrianto and Sharmanska, 2017; Zafar et al., 2017; Berk et al., 2017; Donini et al., 2018; Raff et al., 2018; Aghaei et al., 2019; Prost et al., 2019; Beutel et al., 2019; Grari et al., 2019; Taskesen et al., 2020; Grari et al., 2020; Cho et al., 2020; Chzhen and Schreuder, 2020; Jiang et al., 2020; Lowy et al., 2022) add a regularizer or constraint to the learning objective. Post-processing approaches (Hardt et al., 2016; Pleiss et al., 2017; Alghamdi et al., 2022; Xian et al., 2023; Ţifrea et al., 2024) improve group fairness properties by altering the final decision of the classifier. See the survey paper by Hort et al. (2022) for a more comprehensive literature survey.
Among all these classical approaches, we believe post-processing approaches are the most compatible for the LLM-based classifiers. Having said that, most post-processing approaches require access to demographic labels at test time, which is infeasible, especially for LLM-based classifiers. That is why we only focused on FRAPPÉ (Ţifrea et al., 2024) which works without such access.
Appendix B Risks
There are three risks we would like to call out:
Appendix C LLM-Based Classification Methodology
(a) Zero-Shot
(b) Fine Tuning through Introspection
Figure 2: Classification flow diagrams for zero-shot and fine-tuned LLM-based classification. In both cases, decisions are encouraged via ‘text wrappers’ that nudge the LLM to make a classification decision. In the zero-shot case (Fig. 2(a)) we treat the wrapped text as a prefix and query the LLM for two postfix tokens (such as ‘Yes’ or ‘No’) that represent positive and negative decisions. We apply a softmax to these scores to obtain a probability distribution over the classification result and use this for decision making. In the fine-tuned case (Fig. 2(b)) we assume that the LLM is ‘introspective’ and can supply its activations. We instead query the LLM for the activations of its last layer to serve as an embedding. We collect those embeddings into a design matrix then fit a logistic regression model on that matrix and corresponding labels. The logistic regression model is then used for downstream decision making.
Fig. 2 gives a schematic diagram of the classification methodology described in Sec. 2. Note that we use the following wrapper text for our toxicity classifier: “‘TEXT_TO_CLASSIFY’ Categorize the comment on the basis of toxicity. If the comment is toxic (harmful, malicious, derogatory, threat, insult, identity attack, etc.) output Yes, otherwise output No.”
Appendix D Data and Data Processing
The dataset contains 405,130 training instances, 21,293 validation instances, and 21,577 test instances. We make use of all three splits in our work.
Appendix E Remediation Methodology

### v2

# Inducing Group Fairness in Prompt-Based Language Model Decisions

@@ Abstract
Classifiers are used throughout industry to enforce policies, ranging from the detection of toxic content to age-appropriate content filtering. While these classifiers serve important functions, it is also essential that they are built in ways that minimize unfair biases for users. One such fairness consideration is called group fairness, which desires that different sub-population of users receive equal treatment. This is a well-studied problem in the context of ‘classical’ classifiers. However, the emergence of prompt-based language model (LM) decision making has created new opportunities to solve text-based classification tasks, and the fairness properties of these new classifiers are not yet well understood. Further, the ‘remediation toolkit’ is incomplete for LM-based decision makers and little is understood about how to improve decision maker group fairness while maintaining classifier performance. This work sets out to add more tools to that toolbox. We introduce adaptations of existing effective approaches from the classical classifier fairness to the prompt-based classifier space. We also devise simple methods that take advantage of the new structure of prompt-based decision makers and operate at the prompt level. We compare these approaches empirically on real data. Our results suggest that adaptations of approaches that are effective for classical classifiers remain effective in the LM-based classifier environment. However, there is room for further exploration of prompt-based remediation methods (and other remediation methods that take advantage of LM structure).
Inducing Group Fairness in Prompt-Based Language Model Decisions
James Atwood1  Nino Scherrer2 Preethi Lahoti1Ananth Balashankar1Flavien Prost1Ahmad Beirami1 1Google DeepMind  2Google, Paradigms of Intelligence Team
@@ 1 Introduction
Language models (LMs) have shown impressive performance across many tasks and are now being deployed across high-stakes applications such as financial Wu et al. (2023) or medical Singhal et al. (2023) domains. In particular, zero-shot LM-based classifiers Wei et al. (2022a); Anil et al. (2023) have achieved state-of-the-art performance on several natural language classification benchmarks and are being widely adopted for decision making. More recently, such classifiers are leveraged as a reward signal to align models with AI feedback Bai et al. (2022). Hence, it is important to ask: How fair are the classification decisions made by LMs?
In this paper, we consider two classes of LM-based classifiers: (i) prompted (“out-of-the-box”) LM classifiers and (ii) trained classifiers on top of last-layer embeddings extracted from an LM. We first assess whether these LM-based classifiers satisfy a widely adopted classifier group fairness notion called equal opportunity (EO) Hardt et al. (2016); Prost and Beutel (2020). EO is measured as the difference between the false positive rates (FPR) of different demographic groups, where negative outcome is considered an advantaged class. For example, consider a toxicity detection classifier where being labeled as toxic leads to some content moderation policy. It is therefore desirable for content from all demographics to be falsely marked as toxic with an equal rate. We find that prompted LM classifiers demonstrate a significant gap in FPR across multiple demographic groups in the Civil Comments toxicity detection benchmark  Borkan et al. (2019), with Muslim and Jewish groups having 89% and 48% higher FPR as compared to the Christian group. The gap is further increased when we compare embedding-based classifiers with Muslim and Jewish groups having 124% and 71% more FPR compared to the majority group.
We then benchmark the effectiveness of two types of group fairness remediation techniques: (i) prompting-based and (ii) regularization-based remediation methods. For prompting-based methods, we study the effectiveness of different group-agnostic and group-aware fairness encouraging natural language prompts. In the context of regularization-based methods, we study an post-processing remediation method (Ţifrea et al., 2024) and in-processing method (Prost et al., 2019; Beutel et al., 2019). We find that prompt-based remediation methods are unable to decrease the FPR gap in our experiments - with Muslim and Jewish groups still having FPR about 40% higher than the Christian group.
@@
  • We assess the group fairness of two classes of LM-based classifiers (i.e., prompt-based and embedding-based) and show that they do not satisfy equal opportunity (EO) along identity aspects such as religion, race, ethnicity, sex.
  • We evaluate three different remediation techniques (i.e., prompting, in-processing, and post-processing) within the two studied classes of LM-based classifiers. We find that prompting-based remediations fail to achieve lower false positive rates, and that regularization-based approaches achieve better fairness-performance tradeoffs across both classifier classes.
@@
  • We find that in-processing remediation achieves better fairness-performance trade-offs than post-processing methods, but may not be always a feasible option due to limited access to the internals of the model.
2 Related Work
@@
While LMs have been broadly studied in the generative case for trustworthiness, e.g., diversity, stereotypes, gender bias, and toxicity Nadeem et al. (2021); Liang et al. (2021); Deshpande et al. (2023); Lahoti et al. (2023), their fairness in classification problems remains under-explored. Tamkin et al. (2023) provided a method for evaluating how biased a language model may be by generating hypothetical prompts with group information and making decisions by fitting a mixed effects model. The authors of the Flan-T5 model Chung et al. (2022) published group-level performance of a toxicity classifier fit to the Civil Comments Identity Borkan et al. (2019) dataset. Baldini et al. (2021) explored remediation methods for achieving equalized odds for different embedding-based classifier models. To our knowledge, this is the first paper that proposes and empirically evaluates methods for remediating LM-based classifiers drawn from LLMs with respect to equal opportunity fairness.
There is also rich literature on fairness in classical (i.e. non-LM-based) classification. In this paper we focus on the equal opportunity (EO) notion of group fairness Hardt et al. (2016); Prost and Beutel (2020) which is achieved for a group and classifier when the false positive rate (or false negative rate) of the classifier is the same for instances drawn from that group when compared with instances drawn from the majority group.
Methods for improving group fairness can generally be categorized in three main classes: pre-processing, in-processing, and post-processing methods. Pre-processing algorithms (Feldman et al., 2015; Zemel et al., 2013; Calmon et al., 2017) transform the biased data features to a new space in which the labels and sensitive attributes are statistically independent. In-processing methods (Kamiran et al., 2010; Ristanoski et al., 2013; Quadrianto and Sharmanska, 2017; Zafar et al., 2017; Berk et al., 2017; Donini et al., 2018; Raff et al., 2018; Aghaei et al., 2019; Prost et al., 2019; Beutel et al., 2019; Grari et al., 2019; Taskesen et al., 2020; Grari et al., 2020; Cho et al., 2020; Chzhen and Schreuder, 2020; Jiang et al., 2020; Lowy et al., 2022) add a regularizer or constraint to the learning objective. Post-processing approaches (Hardt et al., 2016; Pleiss et al., 2017; Alghamdi et al., 2022; Xian et al., 2023; Xian and Zhao, 2024; Ţifrea et al., 2024) improve group fairness properties by altering the final decision of the classifier. See the survey paper by Hort et al. (2022) for a more comprehensive literature survey.
Among all these classical approaches, we believe post-processing approaches are the most compatible for the LM-based decision making. Having said that, most post-processing approaches require access to demographic labels at test time, which is infeasible, especially for LM-based classifiers. That is why we only focused on FRAPPÉ (Ţifrea et al., 2024) which works without access to demographic labels.
(a) Prompt-based Classifier
(b) Embedding-based Classifier
Figure 1: Classification flow diagrams for prompt-based and embedding-based classifiers. Decisions are encouraged via ‘text wrappers’ that nudge the LM to make a classification decision. (1(a)) For prompt-based classifiers, we treat the wrapped text as a prefix and query the LM for two postfix tokens (such as ‘Yes’ or ‘No’) that represent positive and negative decisions. We apply a softmax to these scores to obtain a probability distribution over the classification result and use this for decision making. (1(b)) For embedding-based classifiers, we assume that the LM is ‘introspective’ and can supply its activations. We instead query the LM for the activations of its last layer to serve as an embedding. We collect those embeddings into a design matrix then fit a logistic regression model on that matrix and corresponding labels. The logistic regression model is then used for downstream decision making.
@@ 3 Problem Setup
Equal opportunity notion of group fairness.
Consider a binary classifier where the positive outcome is considered to be a disadvantaged class, e.g., being detected as toxic. The widely adopted equal opportunity (EO) notion of fairness requires that different demographics are disadvantaged with the same rate, i.e., the false positive rate (FPR) of the classifier is equal for all demographics (Hardt et al., 2016; Prost and Beutel, 2020; Lowy et al., 2022; Hort et al., 2022). For example, this means that the content from different demographics is desired to be falsely marked as toxic at an equal rate.
LM-based classifiers.
We explore two classes of classifiers: (i) prompt-based LM classifiers, and (ii) trained classifier on top of extracted last-layer embeddings of LMs. The former does not involve any training and leverages the instruction following ability of LMs. In contrast, the latter requires supervised training of a classifier, and is also commonly referred to as a trained linear head on top of the last-layer representation. In both cases, we assume grey-box access to the LM, i.e., access to the logits (for prompt-based classifiers) or last layer activations (for embedding-based classifiers) produced by the language model prior to sampling.
  • Prompt-based LM classifier.
    Fig. 1 gives a schematic diagram of the classification methodology described in Sec. 3. Note that we use the following wrapper text for our toxicity classifier: We prompt an LLM with the toxicity classification task formulated as yes or no question. By assessing the log-probabilities of the “Yes” and “no” tokens and applying a softmax function, we are able to extract a distribution over negative and positive classes. A schematic diagram of the process is given in Fig. 1(a) and our prompt is given below.
    {TEXT_TO_CLASSIFY} Categorize the comment on the basis of toxicity. If the comment is toxic (harmful, malicious, derogatory, threat, insult, identity attack, etc.) output Yes, otherwise output No.
  • Embedding-based classifiers. We prompt an LLM in the exact same way as above, but extract the final-layer activations of an LM for a training set. Using the collected pairs of (activation, label), we then train a logistic regression model to predict the ground truth labels. During inference time, we obtain the last layer embedding of the wrapped classification input (see Figure 1(b)) and feed it into the trained classifier. This approach can be thought of as a form of parameter efficient supervised fine-tuning. A schematic diagram is given in Fig. 1(b) in the appendix.
@@ 4 Experiment Setup
Models.
We use PaLM 2 S Anil et al. (2023) as the base model, and PaLM 2 L in a model transfer experiment.
Dataset & Tasks.
We use the Civil Comments Identity Borkan et al. (2019) dataset as it provides the information necessary to evaluate and remediate both group fairness and performance. For processing details, please see Appendix C.
Evaluation metrics.
@@
5 Remediation Methodology